OpenStudy (anonymous):

the length of a rectangle is more than twice its width, and the area of the rectangle is . Find the dimensions of the rectangle.

OpenStudy (anonymous):

what is the area?

OpenStudy (anonymous):

he length of a rectangle is 1ft more than twice its width, and the area of the rectangle is 66ft^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

okay this is pretty simple. suppose the width is W feet. what can you say about the length?

OpenStudy (anonymous):

ok, so the width would be x, while the length would be 2x+1 so the equation would be 2x^2+x=66 You would factor it to 2x^2+x-66=0 So it would be: (2x-11)(x+6)=0. x=-6,11/2 Since -6 is negative, 11/2 is the width. The length would be 12, since 2(11/2)+1=12. Dimensions= 12 by 11/2.

OpenStudy (anonymous):

what some1 said.

OpenStudy (anonymous):

:)